Use the following information from questions 6-12. In both 2018 and 2019 you expect to receive a W2 for $200,000; $65,000 will be already withheld on your W2 for federal income taxes and your income tax bracket is 32%. The short term capital gains on investments (<1 year) are taxed like income, long term capital gains on investments (>-1 year) are taxed at 15%, and dividends are taxed at 22%. Besides your W2 income, the only income you have to report is from one stock on the 1099-B issued by your broker: On January 1*, 2018, you purchased 100 shares of stock at $50 per share: you based your investment decision on publically available sources which speculate the shares price will increase dramatically at some point over the next two years July 1, 2018, is the ex-dividend date for the stock- that morning you received a special onetime cash dividend of $2 per share and the stock dosed at $48 per share On November 29*, 2018, the company announced great news about a clinical drug trial and the share price begins to rise On November 30*, 2018, you purchased 20 shares of stock at $80 per share On December 1*, 2018, there was a 2 for 1 stock split On December 2nd, 2018, you sold 50 shares of stock at $100 per share . * On January 2, 2019, you sold 70 shares at $90 per share You do not receive any dividends in 2019 On December 31, 2019, you sell any remaining shares at $30 per share All stock prices quoted above are at market close on the day referenced Assume shares that were purchased first were sold first (i.e. FIFO) Assume there are no transaction costs 6. For this question only, assume the investors only goal is to reduce their tax liability in 2018. The vestor waite od se are foc est tax bracket was10% lower. Evaluate the words in underlined italics. True or False? a. True b. False 7. There is a trading rule that goes: Buy the rumor, sell the news! Which of the following actions most likely violates this rule: I. The investor purchased shares on January 1st, 2018 Il. The investor purchased shares on December 1st, 2018 Il The investor sold shares on December 2nd, 2018 IV. This investor sold shares on January 1st, 2019 a. C. d. IV e. V 8. For this question only, assume you did not sell any stock on December 2nd. Your 2018 federal tax refund (liability) would be closest to: a. ($1,044) b. ($1,000) c. ($800) d. $800 e. $956 9. For this question only, assume you sold all of your stock on July 1, 2018. What was your total return on the investment? a. b. c. (8%) (4%) (2%) e. 4%10. After the sale of the stock on December 2nd, 2018, your cost basis per share is closest to: a. $27.50 b. $28.16 c. $29.29 d. $50.83 e. $52.50 11. Your 2018 federal gross income (W2 wages, capital gains, and dividends) would be closest to: a. $203,736 b. $203,825 c. $203,936 d. $203,950 e. $207,450 12. Your 2019 federal tax refund (liability) would be closest to: a. ($456) b. ($245) c. ($244) e. $318
Answer #1

Basic info

tax rates

income tax 32%

short term capital gain 32% (are taxed like income)

longterm capital gain 15%

dividend tax 22%

6.ans is a. TRUE

the investor's goal is to reduce tax. so he will want to hold the shares for more than 1 year to avail the tax bracket of 15% (for long term gain).

because here tax of short-term gain(32%) is higher than longterm gain (15%).

even though the short term tax 10% lower i.e. 32% - 10% = 22% is still higher than 15% (long term tax rate)

7. a. i

one of a trading rule is "buy the rumor sale the news"

it means buy the share according to rumor and sell it on the basis of news break.

on 1st jan 2018 the investor took the decision to buy the shares on the basis of publically available sources which you can say a rumor that is the price is going to be high.

8. ans e 956$

if he didnt sell than the only tax will be dividend tax on dividend i.e. 2 x 100 shares= 200

tax 22% = 44$

toatal tax deducted as per question is 65000

actual tax 64000 for W2 in come and 44$ for dividend

total 64044

refund = 65000-6044= 956$

9. ans d 0%

on 1.7.18

sale price of the stock 48 per share and so there will be loss of 200 if sale is made

50-48= 2

2x 1oo shares= 200$

and the only gain that year is dividend of 200 .

so net gain is 200 - 200 =0

10.ana a.27.5$

ist buy 100 shares @ 50 each = 5000$

2nd buy 20 shares @80= 1600$

total price 6600

after share split the total share will be double i.e 120 x 2= 240

cost per share is 6600/ 240= 27.5$

11. ans a 203736$

12.ans a 318$